OpenStudy (anonymous):

Which of the following equations is quadratic? x(x2 + 1) = 0 5(4x + 2) = 3 (x + 3)(x + 4) = 5

OpenStudy (anonymous):

The third one :)

OpenStudy (anonymous):

And how did you get to that?

OpenStudy (anonymous):

I'll show you the solution in 5 minutes if you don't mind :)

OpenStudy (anonymous):

ok. Sounds good.

OpenStudy (anonymous):

First one is x(x2 + 1) = 0 Solution will be x^3+x=0 Second one is 5(4x + 2) = 3 Solution will be 20x+10=3 Third one is (x + 3)(x + 4) = 5 Solution will be x^2+4x+3x+12=2 Example of Quadratic exquation is x^2+x+3=0 The first and second is linear equation because example of linear equation is x+3=0 I hope I didn't confuse you :)

OpenStudy (anonymous):

I think I mainly understand? But could you explain the solution to the 3rd one? Like how you got x^2+4x+3x+12=2 as the solution?

OpenStudy (anonymous):

Sorry sorry the first is cubic,the second is linear and the third is quadratic

OpenStudy (anonymous):

ohh. Nvm. I understand.

OpenStudy (anonymous):

How do we get (x + 3)(x + 4) = 5 from x^2+4x+3x+12=2 though?

OpenStudy (anonymous):

Umm listen i got it by multiplying both of them sorry the answer is x^2+4x+3x+12=5 so the final answer is x^2+7x+12=5 :)

OpenStudy (anonymous):